### v2

# BanTH: A Multi-label Hate Speech Detection Dataset for Transliterated Bangla

Abstract
The proliferation of transliterated texts in digital spaces has emphasized the need for detecting and classifying hate speech in languages beyond English, particularly in low-resource languages. As online discourse can perpetuate discrimination based on target groups, e.g. gender, religion, and origin, multi-label classification of hateful content can help in comprehending hate motivation and enhance content moderation. While previous efforts have focused on monolingual or binary hate classification tasks, no work has yet addressed the challenge of multi-label hate speech classification in transliterated Bangla. We introduce BanTH, the first multi-label transliterated Bangla hate speech dataset comprising 37.3k samples. The samples are sourced from YouTube comments, where each instance is labeled with one or more target groups, reflecting the regional demographic. We establish novel transformer encoder-based baselines by further pre-training on transliterated Bangla corpus. We also propose a novel translation-based LLM prompting strategy for transliterated text. Experiments reveal our further pre-trained encoders achieving state-of-the-art performance on the BanTH dataset while our translation-based prompting outperforms other strategies in the zero-shot setting. The introduction of BanTH not only fills a critical gap in hate speech research for Bangla but also sets the stage for future exploration into code-mixed and multi-label classification challenges in underrepresented languages.
BanTH: A Multi-label Hate Speech Detection
Dataset for Transliterated Bangla
[…]
Content Warning: This article contains examples of hateful content.
Note: Throughout the work, we use the term Bangla to refer to both Bangla and the endonym Bengali, both denoting the same language primarily spoken by people of the West Bengal region of India and the vast majority of Bangladesh.
1 Introduction
Figure 1: Binary and Multi-label Classification on an example Transliterated Bangla sentence from the BanTH dataset along with its corresponding English translation.
The ever-expanding digital landscape that promises to improve social cohesion has been a breeding ground for hate speech. Hate speech is defined as any form of language that targets, attacks, or incites implicit or explicit forms of hatred or violence against groups, based on specific characteristics, e.g. physical appearance, religion, national or ethnic origin, and gender identity. Hate speech can potentially inflict societal harm by promoting division among communities, exacerbating mental health problems, and inciting violence. The lack of hate speech moderation cultivates an environment of intolerance and magnifies the negative impact on the target communities Hangartner et al. (2021). Furthermore, categorizing hate speech requires a multi-faceted approach to capture overlapping hate categories and gain a granular understanding of the underlying motives behind the hateful discourse.
A prevalent informal form of online communication for non-English languages uses transliterated texts, i.e. writing one language in the script of another. Transliteration is predominant in Bangla, where Latin characters are used to produce a colloquial form of Bangla texts that does not strictly adhere to the original linguistic rules. Despite boasting over a quarter of a billion speakers worldwide, Bangla remains a low-resource language in terms of Natural Language Processing (NLP) tasks. The challenges are compounded when working with the even more under-resourced transliterated Bangla, which dominates online spaces due to the widespread familiarity of users with English keyboard layouts.
Although transliteration facilitates easier typing and cross-linguistic communication, it also complicates the task of automated hate speech detection due to inconsistent spelling and structure, absence of grammar, mixing with English, and the loss of script-specific features Jahan et al. (2019a). While standard practices have employed transformer-based encoders Devlin et al. (2019) for automated hate speech detection, the recent surge in popularity of Large Language Models (LLMs) has positioned them as a viable option in hate speech NLP, particularly in the zero-shot setting. While most of the advancements have been centered on English and other high-resource languages, transliterated Bangla has very limited research on hate or hate-like speech detection Jahan et al. (2019b); Raihan et al. (2023) and LLM-based methods Shibli et al. (2023), with no work addressing the intersection of these areas.
Addressing the aforementioned research gap, our contribution can be summarized as follows:
  • We propose BanTH, the first multi-label hate speech detection dataset on transliterated Bangla with 37,350 samples.
  • We establish several encoder-based baselines along with our novel transliterated Bangla further-pretrained encoders, achieving state-of-the-art accuracy on our dataset.
  • We explore zero-shot and few-shot prompting techniques using state-of-the-art Large Language Models and introduce a novel translation-based prompting strategy that outperforms existing methods on our dataset.
Source Category #Hate #Non-hate
News & Politics 8745 23695
People & Blogs 1125 2889
Entertainment 412 584
Total 10282 27167
Dataset Statistics
Min Character Count 7
Max Character Count 1951
Average Length 58.09
Standard Deviation 65.13
Total Words 379411
Average Words 10.16
Min Word Count 3
Max Word Count 368
Unique Words 75977
BanTH split
- Train 29879
- Val 3736
- Test 3735
Total 37350
Table 1: Dataset statistics of the BanTH dataset. Min Length and Max Length are the minimum and maximum sentence lengths in terms of the number of characters, respectively. The Std Length indicates the standard deviation of the number of characters per sentence.
2 BanTH Dataset
Figure 2: Overview of the BanTH dataset creation pipeline.
The BanTH dataset comprises 37350 samples, each initially classified into binary labels of Hate or Non-Hate. To effectively capture the target group, the hate samples are further multi-labeled as Political, Religious, Gender, Personal Offense, Abusive/Violence, Origin, and Body Shaming. A detailed description of the target group is provided in Appendix C.1. The flowchart of the dataset creation is given in Fig. 2.
2.1 Data Sourcing
We construct the dataset by scraping user comments from public YouTube videos using the YouTube API [^1] https://developers.google.com/youtube/v3/getting-started. The scraped comments include three categories of videos: “News & Politics”, “People & Blogs”, and “Entertainment”, totaling 26 different YouTube channels, outlined in Table 1. The contents often cover a wide range of topics, including political analysis, interviews, talk shows, product reviews, and travel videos, regionally relevant to West Bengal of India and Bangladesh.
The scraped comments were collected from January 2020 to July 2024, a time frame that captures significant recent events, e.g. the COVID-19 pandemic and the Bangladesh quota reform movement [^2] https://en.wikipedia.org/wiki/2024_Bangladesh_quota_reform_movement. Such events led to increased social media activity. The variety of content appeals to a diverse audience with a wide range of interests and perspectives. The label-wise distribution of video categories is elaborated in Appendix C.2.
2.2 Data Filtering
We primarily filtered comments containing English characters. During annotation, non-Bangla and non-English transliterated comments, such as comments in Hindi, were discarded. The scraped comments were filtered through English character filtering and were further filtered using dictionary-based language identification. Original English and Bangla and code-mixed texts were filtered out using the dictionary filtering approach. Transliterated texts in any other language such as Hindi, were dropped during annotation. Only the transliterated Bangla texts with lengths greater than or equal to three were kept in the dataset.
2.3 Data Cleaning
During data cleaning, we removed URLs using regular expressions as they were not relevant to our datasets. Comments that were verbatim duplicates of others were removed. Personal information like addresses and contact numbers were also removed during annotation.
2.4 Data Annotation
We hired four annotators and two domain experts for the data annotation process. All the annotators were high school graduates and native Bangla speakers. They had extensive exposure to social media content and actively used transliterated Bangla text. The annotators were compensated fairly for their work, with a rate of BDT 1 per sample. The domain experts were remunerated hourly. The annotation process was supervised by domain experts with experience in working with transliterated Bangla text, who provided guidance and ensured clear and proper annotation guidelines, outlined in Appendix B.
Each data sample was annotated by three annotators to ensure consistency and capture a shared perspective. Each data sample was first instructed to be classified as hate or non-hate and each resulting hate sample was then instructed to be multi-labeled according to the target group descriptions given above. In cases where annotators encountered difficulties in reaching a consensus or understanding the text, these issues were addressed and instantly resolved by the domain experts to avoid further confusion.
Fleiss’ Kappa Expert- Annotator Inter- Annotator
Binary Labeling 0.75 0.71
Political 0.62 0.72
Religious 0.68 0.66
Gender 0.64 0.68
Personal Offense 0.70 0.75
Abusive/Violence 0.72 0.71
Origin 0.68 0.72
Body Shaming 0.64 0.66
Table 2: Fleiss’ Kappa Score of annotators and domain experts across categories.
2.5 Data Validation
Following the annotation process, the dataset underwent a thorough review by the domain experts to ensure labeling accuracy. The experts verified whether any samples remained unlabeled, whether any samples were mislabeled, or if there were any instances where a sample was labeled as hate without the corresponding multi-label, and vice versa. Additionally, the dataset was examined by the domain experts to ensure that all samples were in transliterated Bangla, as samples in other languages were considered beyond the scope of the study and were subsequently removed. The agreement between the annotators the domain experts, and the inter-annotator was measured using the Fleiss’ Kappa metric and the result is shown the Table 2. An inter-annotator agreement score of 0.53 indicates a moderate agreement across the dataset Islam et al. (2021).
2.6 Dataset Statistics
Following data validation, the dataset is split into train, test, and val datasets maintaining the standard 80:10:10 ratio. During the split, a stratification approach was followed to maintain a uniform distribution of hate and non-hate labels and further multi-label categories across all the splits. Further data statistics are given in Table 1.
Figure 3: Heatmap of co-occurrence of multi-label Hate in the BanTH dataset.
The heatmap in Figure 3 provides a detailed visualization of multi-label co-occurrence within the BanTH dataset, illustrating the frequency with which pairs of labels appear together in the text data. Each cell in the heatmap represents the count of co-occurrence between two specific labels, with darker shades of purple indicating higher levels of co-occurrence. The diagonal cells depict the total occurrences of each label independently, with Personal Offense emerging as the most frequent category, appearing in 6966 samples. Conversely, Body Shaming is the least frequent label in the dataset. Significant co-occurrence patterns are observed between Personal Offense and Abusive/Violence with 618 samples, as well as between Personal Offense and Political with 472 samples. In contrast, labels such as Religious and Body Shaming demonstrate minimal co-occurrence with other categories, reflecting their relative isolation in the dataset. The heatmap underscores the prominence of Personal Offense in the BanTH dataset, indicating that a substantial portion of the text data is classified under this label.
3 Methodology
We use various baseline approaches to assess the quality of BanTH, establishing the benchmark performance of it. The experiments are carried out in two stages: Binary Classification of the dataset, followed by Multi-label classification of the hate speech predicted in the first stage. For evaluation, we follow two approaches: one involves finetuning the pre-trained LMs on our corpus, BanTH, and the other involves prompting state-of-the-art large GPT-based LLMs. The following sections discuss the experiments in detail.
3.1 Further Pretraining
Pretrained LMs such as BERT Devlin et al. (2019) and BanglaBERT Bhattacharjee et al. (2022) are typically trained on general and formal texts. However, these models often lack transliterated samples in their pretraining datasets. To enhance the performance of these models, we employ the technique of Further Pretraining (FPT) Qiu et al. (2021); Gururangan et al. (2020), utilizing the Masked Language Modeling (MLM) loss Devlin et al. (2019); Zhuang et al. (2021) as our pretraining objective. This approach aims to improve the models’ effectiveness on our BanTH dataset.
Specifically, for a given sentence $S$, we represent the input tokens as $X=(x_{1},x_{2},\ldots,x_{n})$. We randomly mask 15% of these tokens with the [MASK] token, resulting in a masked sequence denoted as $\tilde{X}$. Let $M\subseteq\{1,2,\ldots,n\}$ represent the set of indices of the masked tokens. We then input $\tilde{X}$ into the language model $f_{LM}$. For each masked token $x_{i}$ (where $i\in M$), the model generates predicted probabilities $P(x_{i}|\tilde{X})$. The loss for each masked token is calculated as $L_{i}=-\log P(x_{i}|\tilde{X})$. The overall loss across all masked tokens is given by:
$L_{\text{MLM}}=-\frac{1}{|M|}\sum_{i\in M}\log P(x_{i}|\tilde{X})$
For our FPT process, we utilize the BanglaTLit-PT corpus [^3] https://www.kaggle.com/datasets/farihatanjimshifat1/bangla-transliteration-further-pretraining-dataset, which contains 243k unlabeled Bangla transliterated texts. After further pretraining the models on BanglaTLit-PT, we build Bangla transliteration-enhanced encoder models namely TB-Encoders (Transliterated Bangla Encoders).
3.2 LM Finetuning
In our fine-tuning experiments, we utilize a pre-trained language model $f_{\text{LM}}$ that has been specifically trained on our dataset. We input a sentence $\mathbf{S}$ into $f_{\text{LM}}$ to obtain layer-wise contextual representations $\mathbf{H}=\{\mathbf{h}_{i}^{l}\}_{l=1}^{L}$, where $\mathbf{H}=f_{\text{LM}}(S)$ and $L$ = no. of layers in $f_{\text{LM}}$. We use the representation from the last layer’s [CLS] token, $h^{L}_{\text{CLS}}$, as the sentence representation, which is then passed through a Multi-Layer Perceptron (MLP) for classification. The transformation is defined as follows:
$z=W_{2}\cdot(\text{ReLU}(W_{1}\cdot h^{L}_{\text{CLS}}+b_{1}))+b_{2}$
The resulting representation $z$ is then employed for calculating the loss. In the case of TB-Encoders, we also similarly fine-tune those models after doing FPT on the BanglaTLit-PT dataset.
Prompt-Strategy Prompt
Non-Explanatory Base-prompt
Chain of Though (CoT) Base-prompt + "Let’s think step by step"
Explanation-based (Exp) Base-prompt + "Explain why"
HARE Base-prompt + "Let’s explain step by step"
Why [Positive] Base-prompt + "Explain why the comment is positive"
Why [Negative] Base-prompt + "Explain why the comment is negative"
Translation Based [BAN] "Translate the following transliterated text into standard Bangla" + Prompt-Strategy
Translation Based [ENG] "Translate the following transliterated text into standard English" + Prompt-Strategy
Table 3: The prompting strategies used in our benchmarks. We consider binary/multi-label classification and zero-shot/few-shot approaches along with their corresponding prompts.
3.3 Prompting Strategy
We adopt one of the following strategies: Why Positive/Why Negative Wang et al. (2023), HARE Yang et al. (2023), non-explanatory, explanation-based, CoT based or translation based. An overview of the prompts has been reported in Tab. 3. We consider GPT 4o and GPT-3.5 in the prompting experiments and have designed separate base prompts for binary and multi-label classification. Each of these base prompts is then extended to one of the prompting strategies. We consider both zero-shot and few-shot prompting techniques, reported in App. E. Overall, all the variations of the prompts can be categorized as follows:
Non-Explanatory. The non-explanatory prompting strategy forms the base prompts for other prompting strategies, where we only ask the LLM to do binary or multi-label classification. The prompt also includes the hate speech definition, the target groups, and their definition in multi-label classification, key indicators of hate, and labeling instructions.
Explanation Based. In the explanation-based prompting approach, we include the question Explain why with the base prompt and ask the LLM to return its reasoning.
HARE. In HARE-based prompting strategy, we insert the line Let’s explain step by step, as adapted in Yang et al. (2023), along with the base-prompt and asked the LLM to return its reason for the corresponding binary or multi-label.
Why Positive Why Negative. As explained in Wang et al. (2023), we adopt the approach by adding the question why the given text is positive or negative appending to the base prompt in this variation.
CoT-Based. CoT-based prompting strategy is implemented by adding the question Let’s think step by step along with the base prompt.
Translation Based. LLMs are trained primarily on extensive formal text, which may limit their performance and explanatory capabilities when handling transliterated text. To address this issue, we propose a simple yet effective prompting strategy called the Translation-Based prompting strategy. This approach requires LLMs to first convert Bengali transliterated text into standard Bangla or English before proceeding with the usual classification. We test this proposed prompting method using Non-Explanatory, Explanatory, and CoT-based prompting techniques.
All the prompts are provided in Appendix E.
Model Label Macro-F1 $\uparrow$ Acc. $\uparrow$
GPT 3.5 Positive 62.48 69.03
Negative 62.71 66.77
GPT 4o Positive 63.33 74.91
Negative 69.98 75.29
Table 4: Performance of Why prompting on the BanTH test split for binary classification.
Models Binary Multi-label
Macro-F1 $\uparrow$ Acc. $\uparrow$ Macro-F1 $\uparrow$ Subset Acc. $\uparrow$ Hamming Loss $\downarrow$
Language Model (LM) Fine-tuning
Bangla LM
   BanglaBERT 76.50 81.04 20.82 54.08 7.26
   BanglishBERT 75.07 80.62 20.61 52.74 7.42
   BanglaHateBERT 70.92 77.54 11.34 49.83 7.97
   VAC-BERT 74.19 79.76 18.45 52.56 7.50
Indian LM
   MuRIL 75.29 80.83 14.58 53.98 7.55
   IndicBERT 74.51 80.48 13.39 51.56 7.88
Multilingual LM
   mBERT 74.97 80.37 28.24 52.19 7.78
   XLM-R 77.35 81.37 29.29 53.28 7.23
Character-based LM
   CharBERT 76.61 80.91 19.66 53.21 7.44
Transliterated Bangla (TB) LM (Ours)
   TB-BERT 76.27 79.25 30.17 54.19 7.18
   TB-mBERT 77.36 82.57 27.07 54.71 7.28
   TB-XLM-R 77.04 81.29 29.04 52.86 7.26
   TB-BanglaBERT 77.12 81.61 22.52 53.97 7.37
   TB-BanglishBERT 77.12 81.39 21.41 52.79 7.42
Large Language Model (LLM) Prompting
Non-Explanatory
   GPT 3.5 61.44 64.02 24.12 18.27 19.77
   GPT 4o 70.05 74.30 36.07 22.60 16.91
   GPT 3.5 + Few-shot 61.85 65.65 23.86 16.85 20.36
   GPT 4o + Few-shot 68.77 74.18 39.53 26.76 14.16
Chain of Thought (CoT) Prompting
   GPT 3.5 + CoT 61.87 65.77 25.61 20.34 19.28
   GPT 4o + CoT 69.87 74.14 35.97 22.60 16.64
   GPT 3.5 + CoT + Few-shot 63.30 67.80 28.35 20.15 19.15
   GPT 4o + CoT + Few-shot 69.50 74.94 36.49 21.10 17.36
Explanation-based (Exp) Prompting
   GPT 3.5 + Exp 61.69 65.16 22.61 17.33 19.84
   GPT 4o + Exp 69.91 74.16 32.58 19.12 17.94
   GPT 3.5 + Exp + Few-shot 62.69 66.60 25.00 18.36 19.57
   GPT 4o + Exp + Few-shot 70.70 75.15 35.61 21.33 17.15
HARE Prompting
   GPT 3.5 + HARE 62.64 66.77 25.99 20.90 18.94
   GPT 4o + HARE 69.71 74.75 36.67 20.80 17.10
   GPT 3.5 + HARE + Few-shot 62.20 66.72 24.68 20.06 19.18
   GPT 4o + HARE + Few-shot 69.12 74.75 34.06 21.59 17.37
Translation Prompting (Ours)
   GPT 4o + Translation [BAN] 69.63 72.70 35.72 23.28 16.73
   GPT 4o + Translation [ENG] 69.38 72.96 34.28 20.72 16.91
   GPT 4o + Translation [BAN] + CoT 69.01 72.66 36.22 21.35 17.36
   GPT 4o + Translation [ENG] + CoT 69.74 74.87 35.94 21.19 16.91
   GPT 4o + Translation [BAN] + Exp 70.34 73.65 36.20 21.54 16.46
   GPT 4o + Translation [ENG] + Exp 70.19 73.84 36.25 23.92 16.23
Table 5: Model Benchmarking on BanTH Dataset on test split for Binary and Multi-label Classification. For prompting LLMS the results show the zero-shot and few-shot approaches for each of the strategies. All the Translation Prompting is designed following a zero-shot approach. Bold represents overall the best results whereas Underline represents the best performing results in zero-shot settings.
4 Experimental Result
We conduct experiments on fine-tuning LMs along with utilizing prompt-based techniques. Our experiment setup for finetuning LMs is described in Appendix D and the prompts are reported in Appendix E. Table 5 shows the benchmarking of various models on our dataset for both binary and multi-label classification tasks.
4.1 Fine-Tuning LMs Baseline
For the binary classification, among the fine-tuned models, TB-mBERT achieves the highest accuracy (82.57%) and Macro-F1 score (77.36%). The best-performing BERT variant is TB-BanglaBERT, with an accuracy of 81.61% and Macro-F1 of 77.12%. In multi-label classification, for Macro-F1, TB-BERT leads with 30.17%, followed closely by XLM-RoBERTa (28.24%). In subset accuracy, TB-mBERT performs best (54.71%), with TB-BERT close behind (54.19%). Hamming Loss is lowest (best) for TB-BanglaBERT (7.26). From the experiments, our proposed TB-encoders beat the other LMs for both binary and multi-label classification.
4.2 Prompt Based LLM Performance
Among prompt-based LLMs, GPT-4o + Few-shot performed the best for multi-label classification, achieving the highest Macro-F1 (39.53%), Subset Accuracy (26.76%), and lowest Hamming Loss (14.16%). In binary classification, GPT 4o+Exp+Few-shot has the highest Macro-F1 (70.70%). In terms of accuracy, our Translation Prompting approach gains the highest result (74.87%). If we consider zero-shot prompting techniques, the proposed approach consistently outperforms all other prompt variations.
Our extensive experiments reveal that fine-tuned models outperform all prompt-based results for both binary and multi-label classification with a good margin (~8% improvement over prompting). The few-shot prompting strategy enhances model performance compared to the zero-shot approach, although this is not true for all cases. Additionally, GPT-4 consistently outperforms GPT-3.5 across all experiments and prompting variations, yielding improvements of 8-10%. We also find that generating explanations can occasionally boost performance, but it may also lead to worse results in some instances. Our proposed prompting techniques achieve the best performance in zero-shot settings. Overall, our proposed TB encoders demonstrate superior performance, surpassing all fine-tuned LMs and prompt-based results.
Figure 4: Category-wise classification Macro-F1 score in the BanTH dataset with the best-performing model.
5 Error Analysis
Figure 4 illustrates per-label accuracy with the best performing model, TB-mBert. From the fig, it is evident that the classification accuracy for all the labels except Personal Offense label. Figure 3 indicates that most hate speech labels are categorized under Personal Offense. Consequently, the subset accuracy of the multi-label classification, as presented in Table 5, is lower than the per-label accuracy. This suggests that the model struggles to effectively classify samples of Personal Offense label. As illustrated by Fig. 5, the minimum accuracy for binary and multi-label classification is at least 80% and 50%, respectively, with Entertainment videos holding the highest accuracy in both the classifications, showing higher subset accuracy in multi-label classification (79.5%).
Figure 5: Video category-wise class classification accuracy on the BanTH dataset with the best-performing model. We consider Macro-Accuracy for Binary and Subset-Accuracy for Multi-label Classification.
It is evident from Tab. 6 that the number of Personal Offense in Entertainment video type is comparable to that of the other multi-labels of hate, whereas in the other two categories, News & Politics and People & Blogs, the number of Personal Offense hate is noticeably high. The overall improved accuracy in Entertainment category aligns with the result produced in Fig. 4.
The results in Tab. 5 show that our Translation-based models, like GPT 4o Bangla/English Translation with explanations, perform nearly as well as top prompt-based models. As observed in Tab. 7, LLMs can naturally translate Bangla transliterated text into English, even without explicit instructions. However, when asked to translate, the LLM tends to give more literal translations. We also observe that poor translations often lead to incorrect classifications, especially with regional dialects or terms lacking direct English equivalents. English translations generally outperform Bangla ones. When not prompted to translate, the LLM focuses on identifying hate-related terms, which works for explicit hate speech but struggles with subtler forms of offensive content, leading to misclassification.
6 Related Work
Hate Speech.  Natural Language Processing (NLP) tasks on hate speech primarily involve hate speech detection Schmidt and Wiegand (2017) which can be binary Mutanga et al. (2020); Bose and Su (2022), multi-class Yigezu et al. (2023); Hashmi and Yayilgan (2024), or multi-label classification Mollas et al. (2022). Multi-class and multi-label tasks generally categorize hate based on race, religion, gender, origin, and disability. Hate speech detection has also been explored in multi-modal Gomez et al. (2020); Kiela et al. (2020) and multi-lingual Aluru et al. (2020); Ousidhoum et al. (2019) settings.
Bangla Hate Speech.  As a low-resource language, Bangla has limited work on hate speech detection Romim et al. (2022); Das et al. (2022a) but has made significant progress on other hate-related tasks involving abusive content Hussain et al. (2018); Jahan et al. (2022a), cyberbullying Emon et al. (2022), sexism Jahan et al. (2023), and violence Fahim (2023). The multi-label variants are generally categorized based on religion, politics, and gender Sharif et al. (2022). Key areas of multi-labeling tasks include aggression Hossain et al. (2023), cyberbullying Saifuddin et al. (2023), hate speech Shakil (2022), and toxic content Belal et al. (2023).
Transliterated Bangla.   Research on NLP tasks using transliterated or romanized Bangla is notably scarce, with limited works on back-transliteration Shibli et al. (2023), abusive content detection Jahan et al. (2019b); Sazzed (2021), hate speech detection Das et al. (2022a), offensive language identification Raihan et al. (2023), sentiment analysis Hassan et al. (2016), and cyberbullying detection Ahmed et al. (2021). While multi-label classification tasks on transliterated Bangla have been explored as sentiment and emotion analysis tasks Tripto and Ali (2018), there remains a significant research gap on multi-label hate speech classification for transliterated Bangla.
7 Conclusion
Our novel multi-label hate speech dataset, BanTH, addresses a crucial research gap in the domain of transliterated Bangla. Extensive experimentations on both traditional language models and LLMs highlight their usefulness in their unique settings. We envision our novel prompting strategies to be generalized to the processing of transliterated text in other languages. We believe our dataset can serve as a valuable resource in the creation of safer digital platforms for the future.
Limitations
The BanTH dataset contains 27167 non-hate samples out of 37350, i.e. a dummy classifier predicting all samples as non-hate achieves an accuracy of 72.73%. This undermines the performance of a few methods, notably, GPT-3.5 performs worse than this baseline. We also recognize that our work does not cover all possible hate classes and may not adequately represent other regionally relevant hate classes, e.g. hate based on socio-economic conditions.
Ethical Considerations
The use of comments scraped from the YouTube videos complies with the YouTube API’s terms of services [^4] https://developers.google.com/youtube/terms/api-services-terms-of-service. All forms of Personal Identification Information (PII) have been removed from the dataset to prevent privacy violations. We ensured fair distribution of annotation workload. The hired annotators and domain experts were compensated on an hourly basis at a rate above the industry standard.
Appendix A Additional Related Work
Hate-like Tasks.  Hate speech lacks a consistent definition across the literature and has significant overlap with similar detection tasks involving offensive language Pradhan et al. (2020); Ranasinghe and Zampieri (2020), sexism Chiril et al. (2021); Shifat et al. (2024), abusive content Nobata et al. (2016), slang Pei et al. (2019), and cyberbullying Verma et al. (2022); Ahmed et al. (2022). The studies on hate speech also focused on areas, such as explainability Mathew et al. (2021); Fahim et al. , fairness Mostafazadeh Davani et al. (2021), and bias mitigation Xia et al. (2020).
Non-English Hate Speech.  While the majority of the hate speech datasets are in US English Tonneau et al. (2024), the task has been explored in several non-English languages including Arabic Anezi (2022), Hindi Das et al. (2022b), Indonesian Ibrohim and Budi (2019), Korean Kang et al. (2022); Lee et al. (2022), Marathi Patil et al. (2022), and Spanish Plaza-del Arco et al. (2021). Due to the prevalence of English alphabets in digital spaces, some of these works involve transliterated Hettiarachchi et al. (2020) and code-mixed texts Bohra et al. (2018).
Appendix B BanTH Dataset Annotation Guidelines
This document outlines the procedures required for accurate annotation of texts within the BanTH dataset. Annotators are expected to follow each step meticulously to ensure consistency and precision. Your role as an annotator is critical to maintaining the integrity and quality of the dataset.
Binary Classification - Hate/Non-Hate:
Texts containing offensive, abusive, or harmful language targeting individuals or groups based on their identity, characteristics, or affiliations are to be labeled hate. If the text does not include offensive or harmful content it is to be labeled non-hate.
Steps:
  1. 1.
    Review the text in transliterated Bangla.
  2. 2.
    Classify it as either Hate Speech or Non-Hate Speech.
  3. 3.
    If the comment is Non-Hate, stop here. No further labeling is necessary.
Multi-Label Hate Speech Classification:
For texts classified as Hate Speech, assign one or more relevant labels from the following categories. Each text may fit multiple categories, so assign all that apply.
  • Political: Texts that attack, criticize, or promote hate based on political affiliations, views, or ideologies. This includes language dehumanizing political opponents or encouraging harm.
    Example:
    [BN]
    "Oder rajniti desh ta noshto korche."
    [EN] "Their politics are destroying the country."
  • Religious: Texts targeting someone’s religious beliefs or lack thereof, or that promote violence or discrimination based on religion.
    Example:
    [BN]
    "Ei dhormo manushkei bhag kore."
    [EN] "This religion only divides people."
  • Gender: Texts that reinforce harmful gender stereotypes or discriminate based on gender identity or expression. Includes sexist or transphobic remarks.
    Example:
    [BN]
    "Meyera kaj korte pare na, barir kaji kore."
    [EN] "Women can’t work, they only do household chores."
  • Personal Offense: Deeply insulting or offensive remarks directed at an individual, which do not target a specific group but are highly personal.
    Example:
    [BN]
    "Tui ekta boka chele."
    [EN] "You’re a stupid boy."
  • Abusive/Violence: Texts that contain explicit threats of violence or encourage violent acts against individuals or groups.
    Example:
    [BN]
    "Tor matha guriye debo."
    [EN] "I’ll smash your head."
  • Origin: Texts targeting someone based on their nationality, ethnicity, or race. This includes racial slurs or calls for exclusion based on origin.
    Example:
    [BN]
    "Ei desher manush er kono kotha nai."
    [EN] "People from this country don’t matter."
  • Body Shaming: Texts that criticize or mock someone’s physical appearance, such as body size or traits, including physical disabilities or conditions.
    Example:
    [BN]
    "Tor pet eto boro kano?"
    [EN] "Why is your stomach so big?"
Each sample will be annotated by three different annotators. After all of this, domain experts will verify the labels to ensure quality.
Notes: Firstly, carefully read and understand each text before assigning labels. Note that, a sample can receive multiple labels if it fits into more than one category. Please ensure cross-verification by collaborating with fellow annotators. If the problem is not solved even after collaboration then contact domain experts because accuracy is essential for a high-quality dataset.
Appendix C Additional Details of BanTH
C.1 Formal Description of Target Groups
The description of the target group with suitable examples is given in the following:
Political: Statements that marginalize, threaten, or incite violence against individuals or groups based on their political affiliations or ideologies. Includes calls for harming political opponents, dehumanizing metaphors applied to political groups, or the deliberate spread of disinformation intended to foment hatred.
Religious: Expressions targeting individuals or groups due to their religious beliefs, practices, or lack thereof. Includes language that demonizes religious communities, advocates for discrimination against religious groups, or incites violence toward religious institutions or adherents.
Gender: Extends beyond sexist remarks and includes language promoting gender-based violence, discrimination based on gender identity, misogynistic or misandrist content, transphobic speech, and language reinforcing harmful gender stereotypes.
Personal Offense: Deeply insulting or offensive language directed at an individual. Includes mocking tragedies, using insulting names, or making derogatory comments.
Abusive/Violence: Explicit threats of violence, detailed descriptions of harm one wishes to inflict on others, or speech that glorifies or encourages violent acts against individuals or groups.
Origin: Expressions that target individuals or groups based on their national, ethnic, or racial background. Includes racial slurs, promotion of racist/nationalist ideologies, calls for racial segregation or deportation, or speech that attributes negative characteristics to groups.
Body Shaming: Derogatory comments on a person’s physical appearance, often related to weight, shape, or size. Can extend to mocking people with physical disabilities, skin conditions, or other visible physical traits.
C.2 Category-wise Multi-label Distribution
Multi-labels #News & Politics #People & Blogs #Entertainment
Political 1005 174 21
Religious 109 30 0
Gender 151 38 4
Personal Offense 5883 784 297
Abusive/Violence 2491 276 121
Origin 201 11 2
Body Shaming 56 29 1
Table 6: Distribution of multi-label hate speech among the YouTube video categories of the BanTH dataset.
The distribution of multi-labels within the BanTH dataset in Table 6 reveals notable variances in the prevalence of various types of hate speech across video categories, including News & Politics, People & Blogs, and Entertainment. Table 1 shows that the News & Politics video category holds a major share in the BanTH dataset, so it is less surprising that the number of hate samples of each label will be greater in that. The Personal Offense and Body Shaming labels are the most and least prominent in the News & Politics category, with 5883 and 56 samples respectively. In fact, the Personal Offense label is prominent in every video category.
C.3 Wordcloud
From Figure 6 and the corresponding English translations in Figure 7, we can visualize the mostly mentioned words for each label in the BanTH dataset. For Political label, people often referred to Hasina (ex-prime minister of Bangladesh), bnp (a major political party in Bangladesh), league/lig (wing of a political party in Bangladesh), dalal (broker), police (a law enforcement force), etc. As we have covered transliterated Bangla YouTube comments of July 2024, the BanTH dataset includes the public responses related to the political unrest of Bangladesh during that period. In Religious labeled sample texts, people often mention Muslim, Hindu nastik (atheist), Allah (the one and only God in Islam), etc. Gender related hate comments are mostly related to females. In Personal Offense case, commenters address dalal (broker) and police (a law enforcement force) contemptuously using tui/tor. The same case is evident in Abusive/Violence labeled samples because the police showed violence during the student protest of July 2024. As transliterated Bangla is used by mainly Indian and Bangladeshi people, Origin related hate is often directed towards Bangladeshi and Indian people. For Body Shaming classified samples, we see people refer to others as mota (fat) and takla (bald) to body shame and it is directed mostly towards females as apu (sister) is one of the most used words. Although the BanTH dataset is concentrated on these types of hateful words, it may not be the case when the source is different.
Appendix D Experimental Setup
For Language Models (LMs) baseline experiments, we fine-tuned a range of pre-trained LMs to establish baselines for binary and multi-label classification on the BanTH dataset. The models include multilingual architectures like XLM-RoBERTaConneau et al. (2019) and mBERTDevlin et al. (2018), as well as language-specific models such as BanglaBERTBhattacharjee et al. (2022) and BanglishBERTBhattacharjee et al. (2022), which are designed for Bangla and code-mixed transliterated Bangla text, respectively. We also employed BanglaHateBERTJahan et al. (2022b), specifically optimized for detecting hate speech in Bangla, along with IndicBERTKakwani et al. (2020) and MuRILKhanuja et al. (2021), tailored for Indian languages. Additionally, we included models like CharBERTMa et al. (2020), and VAC-BERTBhattacharyya et al. (2023) to capture diverse model architectures, including lightweight models and character-level representations.
The models were fine-tuned with a learning rate of 2e-5, using the Adam optimizer. The maximum sequence length was set to 512 tokens, and training was performed for 5 epochs with a batch size of 32. Early stopping was applied based on the validation loss to prevent overfitting. For prompt-based experiments, we used OpenAI API to get LLM responses for GPT-3.5 and GPT-4o models. For measuring the performance, we consider Macro-F1 & Accuracy metrics for binary classification and Macro-F1, Subset-Accuracy & Hamming Loss for multilabel classification. In our further pretraining experiments, we randomly masked 15% of the tokens and used Masked Language Model loss (MLM loss) as the pretraining objectives. We used a learning rate of 1e-5, with batch_size = 32 and we trained the the TB-Encoder models for 5 epochs on the whole training dataset.
D.1 Performance Metrics
F1-Score.   Harmonic mean of precision and recall, providing a balance between the two. It is particularly useful in cases of imbalanced datasets, as it accounts for both false positives and false negatives. For multi-label classification, the F1-Score is computed per label, and the final score is obtained by averaging across all labels.
Macro-Accuracy.   For binary classification, we report macro-accuracy, which averages the accuracy across both classes (positive and negative), providing a balanced view of the model’s performance across the two categories. This is especially relevant when dealing with class imbalance, as it ensures that both classes are equally represented in the final accuracy measure.
Subset-Accuracy.   Requires an exact match between the predicted and true label sets for each instance, meaning that all labels must be predicted correctly for a prediction to be considered accurate Tsoumakas and Katakis (2009).
Hamming Loss Measures the fraction of labels that are incorrectly predicted, either by being wrongly assigned or missed altogether. This metric is useful for multi-label classification as it accounts for both false positives and negatives across all labels. A lower Hamming Loss value indicates better performance as it signifies fewer labeling mistakes.
Appendix E Prompts for Classification on BanTH dataset
Base Prompt for Binary Classification (0 shot) You are an expert at detecting hate speech in Bangla text (written in Latin script/English letters). Your task is to classify texts as either hate speech (true) or non-hate speech (false). HATE SPEECH DEFINITION: Text that expresses or incites hatred, discrimination, or violence against individuals or groups based on: - Political affiliation - Religion (e.g., Hindu, Muslim, Buddhist, Christian) - Gender - Personal offense - Abusive or violence - Origin - Body Shaming KEY INDICATORS OF HATE SPEECH: 1. Dehumanizing language or comparisons 2. Calls for violence or harm 3. Discriminatory slurs or epithets 4. Stereotyping entire groups negatively 5. Promoting supremacy of one group over others LABELING INSTRUCTIONS: 1. For each text, determine if it contains hate speech indicators 2. If any hate indicators are present, classify as hate speech (true) 3. If no hate indicators are present, classify as non-hate speech (false) SPECIAL CONSIDERATIONS: - Consider local context and cultural references - Be mindful of dialectal variations in spelling - Pay attention to code-mixing (Bangla-English hybrid phrases) - For ambiguous cases, focus on the presence of hate indicators You will receive an array of objects, each containing an ’id’ and ’text’. Analyze each text and provide the required classification.
Base Prompt for Multilabel Classification (0 shot) You are an expert at detecting and categorizing hate speech in Bangla text (written in Latin script/English letters). Your task is to classify texts into multiple categories of hate speech. HATE SPEECH CATEGORIES: 1. POLITICAL: Targeting individuals or groups based on political affiliation or beliefs 2. RELIGIOUS: Targeting individuals or groups based on religion (e.g., Hindu, Muslim, Buddhist, Christian) 3. GENDER: Targeting individuals or groups based on gender identity or expression 4. PERSONAL OFFENSE: Insults or attacks directed at specific individuals 5. ABUSIVE/VIOLENCE: Language that is abusive or incites violence 6. BODY SHAMING: Targeting individuals based on physical appearance or body type 7. ORIGIN: Targeting individuals or groups based on their place of origin, ethnicity, or nationality 8. MISC (Miscellaneous): Other forms of hate speech not covered by the above categories KEY INDICATORS OF HATE SPEECH: 1. Dehumanizing language or comparisons 2. Calls for violence or harm 3. Discriminatory slurs or epithets 4. Stereotyping entire groups negatively 5. Promoting supremacy of one group over others LABELING INSTRUCTIONS: 1. For each text, determine if it contains hate speech indicators for any of the defined categories 2. Assign (True) all relevant hate speech categories to the text 3. If no hate indicators are present, classify as non-hate speech (empty array for categories) SPECIAL CONSIDERATIONS: - Consider local context and cultural references - Be mindful of dialectal variations in spelling - Pay attention to code-mixing (Bangla-English hybrid phrases) - A single text may belong to multiple hate speech categories - Use the MISC category for hate speech that doesn’t fit into other categories You will receive an array of objects, each containing an ’id’ and ’text’. Analyze each text and provide the required classification.
In-context Examples in Few-Shot Prompting for Binary Classification HATE(true)
1. Ei hindu gula ke desh theke ber kore dite hobe, era shobi desher shotru
English Translation : These Hindus need to be thrown out of the country, they are all enemies of the nation.
Reason: Calls for the expulsion of religious groups (Hindus) and labels them all as enemies
2. Chakma der manush e bhabo na, oder shob ke mere fela uchit
English Translation : Don’t even think of the Chakmas as humans, all of them should be killed.
Reason: Dehumanizes ethnic group (Chakma) and calls for violence against them
NON-HATE(False)
1. Ami or political opinion er shathe ekmat hoi na, kintu sheta niye kharap lage na
English Translation: I don’t agree with his political opinion, but it doesn’t bother me.
Reason: Expresses disagreement without hatred or discrimination
2. Amader team abar harlo, ki kharap khelse ora
English Translation: Our team lost again, they played so badly
Reason: Expresses criticism without targeting any protected group
CoT-Based Prompt CHAIN-OF-THOUGHT ANALYSIS STEPS: 1. Initial reading: Read the entire text and identify any potential hate speech elements. 2. Keyword identification: Highlight specific words or phrases that could indicate hate speech. 3. Context analysis: Consider the overall context and tone of the message. 4. Category mapping: Connect the identified elements to specific hate speech categories. 5. Multi-label consideration: Determine if the text fits into multiple categories. 6. Intensity assessment: Evaluate the severity or intensity of the hate speech, if present. 7. Final classification: Summarize the findings and assign relevant categories.
In-context Examples in Few-Shot Prompting for Multi-label Classification HATE(true)
1. Ei hindu neta gula ke desh theke ber kore dite hobe, era shobi desher shotru
English Translation: These Hindu leaders need to be thrown out of the country, they are all enemies of the nation.
Categories: RELIGIOUS, POLITICAL
Reason: Calls for expulsion of religious group (Hindus) and labels them all as enemies
2. Tumi ekta mota beyadob, tomar moto lok der vote deoya uchit na
English Translation:You are a rude fat person, people like you don’t deserve to be voted for.
Categories: BODY SHAMING, POLITICAL, PERSONAL OFFENCE
Reason: Uses body shaming language, suggests denying voting rights, and personally insults the individual
3. Ei mohila neta ra shudhu ghor e thakle e desh er unnoti hobe, oder politics e ashar ki dorkar
English Translation: If these women leaders stayed at home, the country would progress; why do they need to come into politics?
Categories: GENDER, POLITICAL
Reason: Discriminates against women in politics, targeting both gender and political participation
NON-HATE(False)
1. Ami or political opinion er shathe ekmat hoi na, kintu sheta niye kharap lage na
English Translation: I don’t agree with his political opinion, but it doesn’t bother me.
Categories: None
Reason: Expresses disagreement without hatred or discrimination
2. Amader team abar harlo, ki kharap khelse ora
English Translation: Our team lost again, they played so badly
Categories: None
Reason: Expresses criticism without targeting any protected group
Translation-Based Prompt You are an expert at detecting hate speech. You will be given an array of Bangla texts (written in Latin script/English letters). First, convert the Bangla Transliterated text into standard Bangla without losing the semantic meaning, then classify the text as either hate speech (true) or non-hate speech (false). Provide a reason for your classification and provide the translation.
Transliterated Text English Translation Pred. Reason by GPT-4o+E Reason by GPT-4o+E+T Translation by GPT-4o+E+T
world er shob deshei torun shomaj montri mp hoy khali Bangladesh chara eitar karon holo onar moto kichu political peraon jara sharajibon nijei kheye jete chay manusher jonno korte chay na In all the countries of the world, youths become social ministers and MPs, but except for Bangladesh. The reason behind this is a kind of political anarchy those who live their entire lives for themselves not for the welfare of people. NH The text criticizes the entire political generation with derogatory implications. The text criticizes political leaders in Bangladesh for not focusing on youth leadership, but it does not incite hatred or use dehumanizing language. In all countries of the world, young people are made ministers or MPs, except in Bangladesh. The reason for this is that some political persons want to selfishly continue benefiting themselves instead of working for the people.
Tomrao to bikri hoe geco You have been sold too NH The text implies corruption and selling out, suggesting negative stereotyping. The phrase does not express hatred or discrimination towards any specific group or individual. They have also been sold out.
Chup catukar khuni Shut up killer sycophant H The text seems to express frustration but lacks direct hate indicators. The phrase ’chup catukar khuni’ implies killing, which incites violence and hate. Shut up, murderous sycophant.
Ore dhora hok ainer awtai ana hok Let him be caught and brought under justice NH The text calls for arrest and punishment. The text talks about bringing someone under the law without expressing or inciting hate or violence. Let him be caught and brought under the law.
Agey apnader mordhe indian gulare khuje ber koren First, find the Indians among yourselves H The text calls for identifying certain individuals without expressing hate. The text stereotypes and discriminates based on origin, which is a hate indicator. First, find the Indian infiltrators among yourselves.
Myb eta India cal parar meye Maybe she is a girl from some slum in India. H The text derogatorily implies an Indian woman based on location, with a stereotype There is no evident expression of hate speech in identifying someone by a phrase or as a resident of a place. Maybe she is a girl from India Calcutta.
eto jhamelar ki orkar era beche thakle sob kiso vole abaro chori bodmashi korbe What’s the use of so much trouble? If they stay alive, they will forget everything and steal and do mischief again. NH The text discusses obstacles without expressing hate. The text suggests that a group of people, if they remain alive, will engage in theft and misconduct again. It stereotypes and implies inherent criminal behavior. What is the point of all this trouble, if they stay alive, they will forget everything and steal and commit mischief again.
Tora nijerai to political party er pasa chatis. You yourselves are the sycophants of political parties. H The text accuses others of opportunistic political affiliations. The text is a general statement about political affiliation without direct hate speech indicators. It appears accusatory but not incite violence or hatred. You people yourselves are involved with the political party.
Bhai ja korso tai paiso..tmra joto marso oitar hisab k dibe ?? Brother, you got what you deserved… will you give an account for how much you have killed? NH The text discusses the consequences of actions without hate indicators. This text contains implicit threat and justification for physical violence, possibly in a retaliatory context. Brother, you got what you did.. Who’s going to account for how much you hit?
Na bnp na awami league Islami Andolon Bangladesh best. Neither BNP nor Awami League, Islami Andolon Bangladesh is the best. NH The text expresses political preference without any hate indicators. This text promotes political supremacy, suggesting one political movement is superior to others. No BNP, no Awami League, Islami Andolon Bangladesh is the best.
Table 7: Comparison of predictions and reasoning generated by GPT-4o using Explanation (E) and Translation (T) prompting. Green indicates that the prediction (H:Hate and NH:Non-Hate) was correctly labeled by GPT-4o+E+T but incorrectly labeled by GPT-4o+E, and red indicates incorrect prediction by both the models.
(a) Political
(b) Religious
(c) Gender
(d) Personal Offense
(e) Abusive/Violence
(f) Origin
(g) Body Shaming
Figure 6: Word clouds constructed from transliterated Bangla texts in the BanTH dataset for each label.
(a) English Political
(b) Religious(English Translated)
(c) Gender (English Translated)
(d) Personal Offense (English Translated)
(e) Abusive/Violence (English Translated)
(f) Origin (English Translated)
(g) Body Shaming (English Translated)
Figure 7: Word clouds constructed from English translations of the Bangla texts in the BanTH dataset for each label.